OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (3x - 7) units and the length of side ON is (x + 13) units. What value of x will make it a rhombus?

OpenStudy (anonymous):

when x=10

hero (hero):

3x-7= x+13 3x-x = 13+7 2x=20 x = 10

OpenStudy (anonymous):

Rhombuses have sides that are congruent... so set the two expressions equal to each other and solve for x :)

OpenStudy (anonymous):

The quadrilateral is a rhombus when the sides are equal in length. The answer is therefore x=10.